OpenStudy (anonymous):

What is the degree measure of an arc 4 ft. long in a circle of radius 10 ft.?

OpenStudy (anonymous):

\[4ft \pi \]

OpenStudy (mertsj):

The circumference of the circle is what?

OpenStudy (goformit100):

\[\theta =l/r\] now try to solve it.

OpenStudy (anonymous):

C = 2 * pi * r

OpenStudy (anonymous):

pi = 3.14 2 * 3.14 * 10 = what

OpenStudy (goformit100):

62.8

OpenStudy (anonymous):

I got these three to choose from 72° 60° 57.6°

OpenStudy (anonymous):

yeah...62.8 360 * 4 * 3.14 /62.8 = 72 degrees. 72 degrees is the answer

OpenStudy (anonymous):

360 * 4 = 1440 1440 * 3.14 = 4521.6 4521.6 / 62.8 = 72

OpenStudy (anonymous):

Thnx

OpenStudy (anonymous):

np

OpenStudy (mertsj):

\[\frac{arc length}{circumference}=\frac{degree measure}{360}\]